(y/(3y-4))-(2/5)=(3y/(3y+4))find the roots
Answer by lwsshak3(11628)

(y/(3y-4))-(2/5)=(3y/(3y+4))find the roots
LCD:5(3y-4)(3y+4)
5y(3y+4)-2(3y-4)(3y+4)=3y*5(3y-4
5y(3y+4)-2(3y-4)(3y+4)=15y(3y-4)
15y^2+20y-2(9y^2-16)=45y^2-60y
15y^2+20y-18y^2+32=45y^2-60y
48y^2-80y-32=0
3y^2-5y-2=0
(3y+1)(y-2)=0
y=-1/3
or
y=2